# Temporal Test-Time Adaptation with State-Space Models

### v2

@@ B.1 Inference
$\hat{\kappa}=\frac{\bar{r}D-\bar{r}^{3}}{1-\bar{r}^{2}}\quad\text{with}\quad\bar{r}=A_{D}(\hat{\kappa}).$ (32)
B.2 Algorithmic Overview
Algorithm 1 STAD-vMF
1:Input: source model $f_{\theta}$, test batches $\{{\mathbf{X}}_{t}\}_{t=1}^{T}$, sliding window size $s$
2:Initialize: mixing coefficients $\pi_{t,k}\leftarrow\frac{1}{K}\forall t,k$, weights ${\mathbf{W}}_{t}\leftarrow{\mathbf{W}}_{0}\forall t$, transition concentration $\kappa^{trans}$ and emission concentration $\kappa^{ems}$
3:for $t\in\mathcal{T}$ do
4:     Define sliding window $S_{t}=\{\tau\mid\max(1,t-s)\leq\tau\leq t\}$
5:     Get representations ${\mathbf{H}}_{t}\leftarrow f_{\theta}^{L-1}({\mathbf{X}}_{t})$
6:     E-step:
7:     for $\tau\in S_{t}$ do
8:         for $k=1,\dots,K$ do
9:              for $n=1,\dots,N_{t}$ do
10:                  Compute $\mathbb{E}[c_{n,\tau,k}]$ by Eqn. 27 $\triangleright$ Cluster assignments
11:              end for
12:              Compute $\mathbb{E}[{\mathbf{w}}_{\tau,k}]$ by Eqn. 28 $\triangleright$ Prototype dynamics
13:         end for
14:     end for
15:     M-Step:
16:     for $\tau\in S_{t}$ do
17:         Compute $\pi_{\tau}$ by Eqn. 31 $\triangleright$ Label distribution
18:     end for
19:     Compute $\kappa^{\mathrm{ems}}$ by Eqn. 30 (optional)
20:     Compute $\kappa^{\mathrm{trans}}$ by Eqn. 29 (optional)
21:     Predict:
22:     Get predictions $\{y_{t,n}\}_{n=1}^{N_{t}}$ by Eqn. 6
23:end for
24:return Predictions $\{y_{t,n}\}_{n=1}^{N_{t}}$
Appendix C Experimental Details
@@
C.1 Source Architectures
@@
  • ResNet: For EVIS, we follow Zhou et al. (2022) and use their ResNet-18 (He et al., 2016) model with a representation dimension of 512 and train on three random seeds. For ImageNet, we follow Song et al. (2023) and employ the standard pre-trained ResNet-50 model from RobustBench (Croce et al., 2021). Latent representations are of dimension 2048.
C.2 Baselines
@@
  • Continual Test-Time Adaptation (CoTTA) (Wang et al., 2022): takes a different approach by optimizing all model parameters with an entropy objective on augmentation averaged predictions and combines it with stochastic weight restore to prevent catastrophic forgetting.
@@
C.3 Implementation Details and Hyperparameters
By the nature of test-time adaptation, choosing hyperparameters is tricky (Zhao et al., 2023b) since one cannot assume access to a validation set of the test distribution in practise. To ensure we report the optimal performance on new or barely used datasets (Yearbook, EVIS, FMoW, CIFAR-10.1 and ImageNetV2), we perform a grid search over hyperparameters as suggested in the original papers. We perform separate grid searches for the uniform label distribution and online imbalanced label distribution setting. Reported performance correspond to the best setting. If the baselines were studied in the gradual CIFAR-10-C setting by Wang et al. (2022), we use their hyperparameter setup; otherwise, we conduct a grid search as described earlier. Unless there is a built-in reset (SAR) or convergence criteria (LAME) all methods run without reset and one optimization step is performed. We use the same batch sizes for all baselines. For Yearbook we comprise all samples of a year in one batch resulting in a batch size of 2048. To create online class imbalance, we reduce the batch size to 64. We use a batch size of 100 for EVIS, CIFAR.10.1 and CIFAR-10-C and 64 for FMoW-Time and ImageNetV2.
@@ C.3 Implementation Details and Hyperparameters
(Wang et al., 2021) Like in BN, the normalization statistics are based on both training and test set. As in Wang et al. (2021), we use the same optimizer settings for test-time adaptation as used for training, except for the learning rate that we find via grid search on $\{1e^{-3},1e^{-4},1e^{-5},1e^{-6},1e^{-7}\}$. Adam optimizer (Kingma & Ba, 2015) is used. For CIFAR-10-C, we follow the hyperparameter setup of Wang et al. (2022) and use Adam optimizer with learning rate $31e-$ 3.
@@
(Wang et al., 2022) We use the same optimizer as used during training (Adam optimizer Kingma & Ba (2015)). For hyperparameter optimization we follow the parameter suggestions by Wang et al. (2022) and conduct a grid search for the learning rate ($\{1e^{-3},1e^{-4},1e^{-5},1e^{-6},1e^{-7}\}$), EMA factor ($\{0.99,0.999,0.9999\}$) and restoration factor ($\{0,0.001,0.01,0.1\}$). Following Wang et al. (2022), we determine the augmentation confidence threshold by the 5% percentile of the softmax prediction confidence from the source model on the source images. For the well-studied CIFAR-10-C dataset, we follow the setting of Wang et al. (2022) and use Adam optimizer with learning rate $31e-$ 3. The EMA factor is set to $0.999$, the restoration factor is $0.01$ and the augmentation confidence threshold is $0.92$.
@@
(Liang et al., 2020) We perform a grid search for the learning rate over $\{1e-3,1e-4\}$ and for $\beta$, the scaling factor for the loss terms, over $\{0.1,0.3\}$.
@@
(Niu et al., 2023) We conduct a grid search over the learning rate selecting among $\{1e-2,1e-3,1e-4,0.00025\}$. Like the authors, we compute the $E_{0}$ threshold as a function of number of classes $0.4\times\ln K$, use SDG, a moving average factor of 0.9, and the reset threshold of 0.2. The updated layers include all of batch, group, and layer normalization where present in the source model.
@@
The hyperparameters are the initialization values of the transition concentration parameter $\kappa^{trans}$, emission concentration parameter $\kappa^{ems}$ and the sliding window size $s$. We chose the concentration parameters from $\{100,1000\}$. Tab. 6 lists employed settings. We use a default window size of $s=3$. For Yearbook, we employ class specific noise parameters $\kappa_{k}^{\mathrm{trans}}$ and $\kappa_{k}^{\mathrm{ems}}$ as discussed in Sec. 3.3. For the other datasets, we found a more restricted noise model beneficial. Particularly, we use global concentration parameters, $\kappa^{\mathrm{trans}}$ and $\kappa^{\mathrm{ems}}$, and follow suggestions by Gopal & Yang (2014) to keep noise concentration parameters fixed instead of learning them via maximum likelihood (see line 17 and 18 in Algorithm 1). Keeping them fix acts as a regularization term as it controls the size of the cluster (via $\kappa^{\mathrm{ems}}$) and the movement of the prototypes (via $\kappa^{\mathrm{trans}}$). Low concentration values generally correspond to more adaptation flexibility while larger values results in a more conservative and rigid model.
@@
Appendix D Additional Results
D.1 Single Sample Adaptation
Table 7: Adaptation accuracy on temporal distribution shift with single sample adaptation (batch size 1) for both covariate shift and additional online label shift: Table shows values as plotted in Fig. 5. Most methods collapse when only provided with one sample per adaptation step. STAD can improve upon the source model in 4 out of 6 scenarios.
@@ D.1 Single Sample Adaptation
D.2 STAD in Combination with BN
@@
D.3 Cluster Visualization

### v3

@@ C.2 Details on STAD-vMF
[…]
Algorithm 3 EM for STAD-vMF
1:Input: prototypes, cluster assignments, mixing coefficients, representations $\{{\mathbf{W}}_{\tau},{\mathbf{C}}_{\tau},\bm{\pi}_{\tau},{\mathbf{H}}_{\tau}\}_{\tau\in S_{t}}$, transition and emission parameters $\kappa^{\mathrm{trans}},\kappa^{\mathrm{ems}}$
2:for $\tau\in S_{t}$ do
3:  E-Step:
4:  Compute: $\lambda_{\tau,n,k}=\frac{\pi_{\tau,k}C_{D}(\kappa^{\mathrm{ems}})\exp\big\{\kappa^{\mathrm{ems}}\mathbb{E}_{q}[{\mathbf{w}}_{\tau,k}]^{\top}{\mathbf{h}}_{\tau,n}\big\}}{\sum_{j=1}^{K}\pi_{\tau,j}C_{D}(\kappa^{\mathrm{ems}})\exp\big\{\kappa^{\mathrm{ems}}\mathbb{E}_{q}[{\mathbf{w}}_{\tau,j}]^{\top}{\mathbf{h}}_{\tau,n}\big\}}$
5:  Get expectation of cluster assignments: $\mathbb{E}_{q}[c_{\tau,n,k}]=\lambda_{\tau,n,k}$
6:  Compute: $\beta_{\tau,k}=\kappa^{\mathrm{trans}}\,\mathbb{I}_{\{\tau>1\}}\,\mathbb{E}_{q}[{\mathbf{w}}_{\tau-1,k}]+\kappa^{\mathrm{ems}}\sum_{n=1}^{N_{\tau}}\mathbb{E}_{q}[c_{\tau,n,k}]{\mathbf{h}}_{\tau,n}+\kappa^{\mathrm{trans}}\,\mathbb{I}_{\{\tau<T\}}\,\mathbb{E}_{q}[{\mathbf{w}}_{\tau+1,k}]$
7:  Compute: $\gamma_{\tau,k}=\lVert\beta_{\tau,k}\rVert$
8:  Compute: $\bm{\rho}_{\tau,k}=\beta_{\tau,k}/\gamma_{\tau,k}$
9:  Get expectation of prototypes: $\mathbb{E}_{q}[{\mathbf{w}}_{\tau,k}]=A_{D}(\gamma_{\tau,k})\bm{\rho}_{\tau,k}$
10:  M-Step:
11:  Compute: $\pi_{\tau,k}=\dfrac{\sum_{n=1}^{N_{\tau}}\mathbb{E}_{q}[c_{\tau,n,k}]}{N_{\tau}}$
12:end for
13:Compute: $\bar{r}^{\mathrm{trans}}=\left\lVert\dfrac{\sum_{\tau\in S_{t}}\sum_{k=1}^{K}\mathbb{E}_{q}[{\mathbf{w}}_{\tau-1,k}]^{\top}\mathbb{E}_{q}[{\mathbf{w}}_{\tau,k}]}{(|S_{t}|-1)\times K}\right\rVert$
14:Compute: $\kappa^{\mathrm{trans}}=\dfrac{\bar{r}^{\mathrm{trans}}D-(\bar{r}^{\mathrm{trans}})^{3}}{1-(\bar{r}^{\mathrm{trans}})^{2}}$
15:Compute: $\bar{r}^{\mathrm{ems}}=\left\lVert\dfrac{\sum_{\tau\in S_{t}}\sum_{k=1}^{K}\sum_{n=1}^{N_{\tau}}\mathbb{E}_{q}[c_{\tau,n,k}]\mathbb{E}_{q}[{\mathbf{w}}_{\tau,k}]^{\top}{\mathbf{h}}_{\tau,n}}{\sum_{\tau\in S_{t}}N_{\tau}}\right\rVert$
16:Compute: $\kappa^{\mathrm{ems}}=\dfrac{\bar{r}^{\mathrm{ems}}D-(\bar{r}^{\mathrm{ems}})^{3}}{1-(\bar{r}^{\mathrm{ems}})^{2}}$
17:Assign: ${\mathbf{W}}_{t}=(\bm{\rho}_{t,1},\dots,\bm{\rho}_{t,K})$
18:Assign: ${\mathbf{C}}_{t}=(\bm{\lambda}_{t,1},\dots,{\bm{\lambda}}_{t,N_{t}})$
19:return ${\mathbf{W}}_{t},{\mathbf{C}}_{t},\bm{\pi}_{t},\kappa^{\mathrm{trans}},\kappa^{\mathrm{ems}}$
Appendix D Experimental Details
@@
D.1 Datasets
  • EVIS: the evolving image search (EVIS) dataset (Zhou et al., 2022b) consists of images of 10 electronic product and vehicle categories retrieved from Google search, indexed by upload date. The dataset captures shift caused by rapid technological advancements, leading to evolving designs across time. It includes 57,600 RGB images of 256x256 pixels from 2009 to 2020. Models are trained on images from 2009-2011 and evaluated on images from 2012-2020.
D.2 Source Architectures
@@
  • ResNet: For EVIS, we follow Zhou et al. (2022b) and use their ResNet-18 (He et al., 2016) model with a representation dimension of 512 and train on three random seeds. For ImageNet-V2, we follow Song et al. (2023) and employ the standard pre-trained ResNet-50 model from RobustBench (Croce et al., 2021). Latent representations are of dimension 2048.
  • ViT-base: For ImageNet-C, we employ ViT-base (Dosovitskiy et al., 2021) from the timm library (Wightman, 2019). Latent representations are of dimension 768.
D.3 Baselines
@@
  • Continual Test-Time Adaptation (CoTTA) (Wang et al., 2022b): optimizes all model parameters with an entropy objective on augmentation averaged predictions and combines it with stochastic weight restore to prevent catastrophic forgetting.
@@
  • Robust Test-time Adaptation (RoTTA) (Yuan et al., 2023) proposes a robust BN layer and the use of a class-balanced memory bank to address simultaneous covariate and label shift.
@@
D.4 Implementation Details and Hyperparameters
By the nature of test-time adaptation, choosing hyperparameters is tricky (Zhao et al., 2023b) since one cannot assume access to a validation set of the test distribution in practise. To ensure we report the optimal performance on new or barely used datasets (Yearbook, EVIS, FMoW, CIFAR-10.1 and ImageNetV2), we perform a grid search over hyperparameters as suggested in the original papers. We perform separate grid searches for the uniform label distribution and online imbalanced label distribution setting. Reported performance correspond to the best setting. If the baselines were studied in the gradual CIFAR-10-C setting by Wang et al. (2022b), we use their hyperparameter setup; otherwise, we conduct a grid search as described earlier. Unless there is a built-in reset (SAR) or convergence criteria (LAME) all methods run without reset and one optimization step is performed. We use the same batch sizes for all baselines. For Yearbook we comprise all samples of a year in one batch resulting in a batch size of 2048. To create online class imbalance, we reduce the batch size to 64. We use a batch size of 100 for EVIS, CIFAR.10.1 and CIFAR-10-C and 64 for FMoW-Time and ImageNetV2.
@@ D.4 Implementation Details and Hyperparameters
(Wang et al., 2021) Like in BN, the normalization statistics are based on both training and test set. As in Wang et al. (2021), we use the same optimizer settings for test-time adaptation as used for training, except for the learning rate that we find via grid search on $\{1e^{-3},1e^{-4},1e^{-5},1e^{-6},1e^{-7}\}$. Adam optimizer (Kingma & Ba, 2015) is used. For CIFAR-10-C, we follow the hyperparameter setup of Wang et al. (2022b) and use Adam optimizer with learning rate 1e31e-3.
@@
(Wang et al., 2022b) We use the same optimizer as used during training (Adam optimizer Kingma & Ba (2015)). For hyperparameter optimization we follow the parameter suggestions by Wang et al. (2022b) and conduct a grid search for the learning rate ($\{1e^{-3},1e^{-4},1e^{-5},1e^{-6},1e^{-7}\}$), EMA factor ($\{0.99,0.999,0.9999\}$) and restoration factor ($\{0,0.001,0.01,0.1\}$). Following Wang et al. (2022b), we determine the augmentation confidence threshold by the 5% percentile of the softmax prediction confidence from the source model on the source images. For the well-studied CIFAR-10-C dataset, we follow the setting of Wang et al. (2022b) and use Adam optimizer with learning rate 1e31e-3. The EMA factor is set to $0.999$, the restoration factor is $0.01$ and the augmentation confidence threshold is $0.92$.
@@
(Liang et al., 2020) We perform a grid search for the learning rate over $\{1e^{-3},1e^{-4}\}$ and for $\beta$, the scaling factor for the loss terms, over $\{0.1,0.3\}$.
@@
(Niu et al., 2023) We conduct a grid search over the learning rate selecting among $\{1e^{-2},1e^{-3},1e^{-4},0.00025\}$. Like the authors, we compute the $E_{0}$ threshold as a function of number of classes $0.4\times\ln K$, use SDG, a moving average factor of 0.9, and the reset threshold of 0.2. While (Niu et al., 2023) apply SAR only to models with layer or group normalization and update those layers, we also evaluate SAR on source architectures with batch normalization, following prior work (Zhao et al., 2023b). As noted in (Niu et al., 2023), BN layers are less effective for small batch sizes, which accounts for the reduced performance of SAR in this setting (Figs.˜4 and 10).
RoTTA
(Yuan et al., 2023) We conduct a search over the learning rate among $\{1e^{-3},1e^{-4},1e^{-5}\}$. As in the experiments of the original paper, we use the default values for $\lambda_{t}=1.0$, $\lambda_{u}=1.0$, $\alpha=0.05$ and $\nu=0.001$.
@@
The hyperparameters are the initialization values of the transition concentration parameter $\kappa^{trans}$, emission concentration parameter $\kappa^{ems}$ and the sliding window size $s$. We chose the concentration parameters from $\{100,1000\}$. Tab.˜7 lists employed settings. We use a default window size of $s=3$. For Yearbook, we employ class specific noise parameters $\kappa_{k}^{\mathrm{trans}}$ and $\kappa_{k}^{\mathrm{ems}}$ as discussed in Sec.˜3.3. For the other datasets, we found a more restricted noise model beneficial. Particularly, we use global concentration parameters, $\kappa^{\mathrm{trans}}$ and $\kappa^{\mathrm{ems}}$, and follow suggestions by Gopal & Yang (2014) to keep noise concentration parameters fixed instead of learning them via maximum likelihood (line 13 - 16 in Algorithm˜3). Keeping them fix acts as a regularization term as it controls the size of the cluster (via $\kappa^{\mathrm{ems}}$) and the movement of the prototypes (via $\kappa^{\mathrm{trans}}$). Low concentration values generally correspond to more adaptation flexibility while larger values results in a more conservative and rigid model.
@@
Appendix E Additional Results
Table 8: Accuracy on domain adaptation benchmarks under covariate shift and uniform label distribution. The source model is trained on three domains and tested on the remaining one, rotating the test domain for averaging.
Method PACS
Source model 82.99 $\pm$ 8.87
adapt feature extractor
BN 82.85 $\pm$ 9.57
TENT 85.30 $\pm$ 7.33
CoTTA 83.59 $\pm$ 8.46
SHOT 83.30 $\pm$ 9.01
SAR 85.03 $\pm$ 7.71
RoTTA 85.11 $\pm$ 7.73
adapt classifier
LAME 83.31 $\pm$ 8.90
T3A 83.68 $\pm$ 9.14
STAD-vMF (ours) 83.91 $\pm$ 8.58
E.1 Domain Adaptation Benchmarks
To study the limitations and applicability of our method STAD, we also test adaptation performance on non-gradual shifts. For that, we use the domain adaptation benchmark PACS, which comprises images of 10 classes across four categorical domains (photo, art-painting, cartoon and sketch). We use DomainBed (Gulrajani & Lopez-Paz, 2021) to train a ResNet-50 model with BN. We test two settings. In the first setting, we follow Iwasawa & Matsuo (2021); Jang et al. (2023) and train the model on three domains and adapt it on the held-out domain. In the second setting, we follow Gui et al. (2024) and train the model on the photo domain and adapt it on the remaining domains sequentially. The second setting is more difficult as the model is exposed to only a single domain during training and needs to leverage several abrupt changes in the distribution over a longer test stream.
Tab.˜8 shows adaptation performance for the first setting. Adaptation gains are generally smaller than on corruption datasets, with TENT, RoTTA, and SAR improving most upon the source model by 2–3 percentage points. While STAD achieves a more modest improvement of 1 percentage point, it remains the best-performing method among classifier adaptation approaches.
Results for the second setting are shown in Tab.˜9. Consistent with Gui et al. (2024), we observe a decreasing performance across all TTA methods over the course of adaptation, highlighting the challenge posed by multiple non-gradual domain shifts. Nevertheless, all TTA methods, except LAME, improve upon the source model by over 10 percentage points on average. Despite the highly non-gradual nature of this test setting, STAD-vMF performs comparably to the baselines, achieving the third-best performance overall. These findings strengthens the results in Sec.˜5.2, which suggest that STAD is applicable beyond gradual, temporal distribution shifts similar as other TTA methods.
Table 9: Accuracy on domain adaptation benchmarks under covariate shift and uniform label distribution. The source model is trained on the photo domain and TTA methods adapt to the remaining domains sequentially. Results show average over three random training seeds. N/A indicates that adaptation is not applied to the source domain.
Domain
Method P $\rightarrow$ A $\rightarrow$ C $\rightarrow$ S Mean
Source 99.34 $\pm$ 0.57 63.10 $\pm$ 1.55 38.37 $\pm$ 4.99 41.51 $\pm$ 2.99 47.66 $\pm$ 1.29
adapt feature extractor
BN N/A 68.03 $\pm$ 1.98 61.15 $\pm$ 0.34 49.64 $\pm$ 0.28 59.61 $\pm$ 0.52
TENT N/A 68.05 $\pm$ 2.11 61.53 $\pm$ 0.52 51.33 $\pm$ 1.35 60.30 $\pm$ 0.28
CoTTA N/A 63.82 $\pm$ 3.23 59.36 $\pm$ 1.35 56.74 $\pm$ 2.70 59.97 $\pm$ 2.17
SHOT N/A 67.91 $\pm$ 2.04 63.17 $\pm$ 0.86 57.90 $\pm$ 1.22 62.99 $\pm$ 0.58
SAR N/A 68.34 $\pm$ 1.81 61.49 $\pm$ 0.36 52.06 $\pm$ 1.10 60.63 $\pm$ 0.67
RoTTA N/A 68.73 $\pm$ 1.13 58.52 $\pm$ 1.32 52.43 $\pm$ 1.19 59.89 $\pm$ 0.19
adapt classifier
LAME N/A 62.73 $\pm$ 1.72 37.80 $\pm$ 5.09 41.01 $\pm$ 2.90 47.18 $\pm$ 1.26
T3A N/A 68.28 $\pm$ 1.55 62.05 $\pm$ 0.67 54.80 $\pm$ 0.97 61.71 $\pm$ 0.40
STAD-vMF (ours) N/A 68.59 $\pm$ 2.51 61.65 $\pm$ 0.39 52.16 $\pm$ 0.49 60.80 $\pm$ 0.72
E.2 Single Sample Adaptation
Table 10: Adaptation accuracy on temporal distribution shift with single sample adaptation (batch size 1) for both covariate shift and additional online label shift: Table shows values as plotted in Fig.˜4. Most methods collapse when only provided with one sample per adaptation step. STAD can improve upon the source model in 4 out of 6 scenarios.
@@ E.2 Single Sample Adaptation
E.3 STAD in Combination with BN
@@
E.4 ImageNet-C
Table 12: Accuracy on gradually increasing corruptions of ImageNet-C using ViT-Base as the source model. BN and RoTTA are specific to batch normalization layers and are thus ineffective on ViT; they are omitted from this experiment. As discussed in Sec.˜5.2, methods that adapt the feature extractor achieve greater adaptation gains on synthetic corruptions—except for CoTTA, which collapses. On this dataset, STAD provides only marginal improvement over the source model.
Corruption severity
Method 1 2 3 4 5 Mean
Source 66.66 59.71 53.88 43.97 32.43 51.33
adapt feature extractor
TENT 68.41 66.01 63.85 59.18 52.89 62.07
CoTTA 48.02 34.87 28.12 19.78 12.71 28.70
SHOT 67.51 64.41 62.33 57.52 51.19 60.59
SAR 67.69 64.70 62.33 57.55 51.01 60.66
adapt classifier
LAME 66.48 59.52 53.64 43.76 32.21 51.12
T3A 66.71 61.31 56.59 46.75 34.22 53.12
STAD-vMF (ours) 66.60 59.81 54.05 44.25 32.80 51.50
E.5 Comparison to Supervised Oracle
We investigate how far STAD, which operates unsupervised and does not require labels, can close the gap to a supervised approach that makes use of labels. Obviously, a supervised approach is superior, as it can directly learn the function mapping input samples to target labels. In contrast, TTA methods rely solely on signals from the input and therefore have strictly less information available. To evaluate how far this gap can be bridged, we continuously fine-tune the source model at each timestep using a small portion of labeled samples. For this, we use another held-out set from the Wild-Time pipeline, which is 10% the size of the adaptation test stream. At each timestep, we fine-tune the model for one epoch on this split and then evaluate the performance on the regular test set.
We find that the supervised model achieves an average accuracy of 90.67% over the entire test stream. Comparing this to the source model at 81.30%, STAD-Gauss at 86.22%, and STAD-vMF at 85.50% (see Tab.˜2), we observe that STAD can partially close the gap between the unadapted source model and the fine-tuned model. Fig.˜7 further reveals that STAD-vMF is on par with the fine-tuned classifier at certain time steps (1986, 1995, 1997). Additionally, we observe that the severity of the temporal distribution shift hampers also the supervised model’s ability to regain in-distribution accuracy. For instance, in the 1970s and 1980s, the performance of the supervised model is 20 points lower than its in-distribution accuracy (nearly 100%).
Figure 7: Accuracy over time for the temporal distribution shift on Yearbook averaged over three random training seeds.
E.6 Cluster Visualization